OpenStudy (zuha):

how to solve: 4x+3+2x-2+6x

OpenStudy (anonymous):

add like terms

OpenStudy (zuha):

you mean: 4x+2x+6x

OpenStudy (zuha):

which will be 12X then should I subtract: -2

OpenStudy (anonymous):

yes. then subtract two from three

OpenStudy (zuha):

so will it be 12x-1= 11x ?!

OpenStudy (zuha):

AM i RIGHT ??

OpenStudy (anonymous):

YES YOU RAE :DDDDDDDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

*ARE

OpenStudy (anonymous):

4x + 3 + 2x - 2 + 6x combine like terms 4x + 2x + 6x = 12x 3 - 2 = 1 so your answer is : 12x - 1 (you cannot subtract 1 from 12 x because they are not like terms.